OpenStudy (anonymous):

Devin wants to have equal amounts of room on all four sides of his ping pong table. A 5ft by 9ft table is placed in the center of a rectangular room that has an area of 221 square feet. What is the width of the clear space Devin will have on all sides of his table?

OpenStudy (turingtest):

The total area is A=Lw=221 sq. ft. Here is the figureIt look like L=5+2x and w=9+2x. So what does our expression for area become?